OpenStudy (anonymous):

solve -7 less than or equal to 1 minus 2x<7 .

OpenStudy (anonymous):

\[-7\leq 1-2x<7\]subtract 1 \[-8\leq -2x<6\] divide by -2, change the inequality \[4\geq x >-3\] is your answer. you can also write \[-3<x\leq 4\]